OpenStudy (goformit100):

Solve this puzzle !! If: 212 = 6, 327 = 9, 624 = 4, 330 = 10, Then: 515 = ?? Reply must If u r a genius..

OpenStudy (goformit100):

@oldrin.bataku

OpenStudy (nurali):

12/2=6, 27/3=9, 24/6=4, 30/3=10, 15/5=3 answer=3

OpenStudy (goformit100):

Thanks

OpenStudy (nurali):

Anytime.